OpenStudy (anonymous):

A triangle has a leg that measures 3 and a hypotenuse of 9. Which leg measure would form a right triangle

OpenStudy (anonymous):

if it's a right triangle, then use Pythagorean theorem to get 9^2=3^2+x^2 do you have answer choices?

OpenStudy (anonymous):

yeah, they are , 12, 90, 72 sqare root, & 90 square root

OpenStudy (anonymous):

well it's 81=9+x^2 so 72=x^2 so which one do you think is the answer?

OpenStudy (anonymous):

i have no idea.

OpenStudy (anonymous):

take the sqrt of both sides of my equation and you get sqrt 72

OpenStudy (anonymous):

oh, thank you.